OpenStudy (anonymous):

where do I start with these? Given the [problem 2x^3-x^2-13x+10 has a quotient of x^2-3x+1 and a remainder of 5, which of the following statements it Not true?

OpenStudy (anonymous):

A. factor of 2x^3-x^2-`13x+10 b. 2x^3-x^2-13x+10= (x^2-3x+1)(2x+5)+5 c.2x^3-x^2-13x+10/2x+5=x^2-3x+1+5/2x+5